OpenStudy (anonymous):

write the expression 3x^12+20x^6+6 in quadratic form

OpenStudy (mr.math):

\[3(x^6)^2+20(x^6)+6\] You can write \(u=x^6\).

OpenStudy (anonymous):

Treat x^6 = y The equation will then be: 3y^2+20y+6